OpenStudy (anonymous):

is (-4,2) a solution of 2x-5y=10

OpenStudy (solomonzelman):

plug it in 2(-4)-5(2)=10 -8-10=10 -18=10 this is not true therefore (-4,2) is not a solution!

OpenStudy (solomonzelman):

Do you see how I did this? Any more questions? I am more than happy to help you out!